Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Thursday, November 5, 1998

The following is a brief summary of the environmental monitoring activities that were completed on November 5, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

11/05/98 (Thursday)

Segment 1 (Emidio to ANF): Report for 11/05/98. Dewatering of the hydrotest for the pipe segment from Copco rd (EMP 33.34) to Peace Valley road (EMP 25.46) occurred last night. Headers were cut from the pipe and preparations were made to continue testing the segment. Trench excavation continued along Grapevine rd ( from south to north) to approximately EMP 16.25. Exclusion fencing and silt fencing was installed along the edge of the road/ROW near this location. Air compressors were staged at Lebec road, EMP 25.19 for dewatering the directional drill segment there. Trench was excavated on the Tejon Bench at EMP 18.45. Pipe was strung near EMP 17.55. Electrical work and construction work continued at Emidio station and the Grapevine pump station and substation. Bore equipment was delivered to the bore site for Grapevine creek, EMP 15.35. Shading and compaction were completed from EMP 6.75 to EMP 7.5, approximately. Backfilling took place along Hwy. 166, where water from the trench was pumped into a water truck for removal, a small hoe opened trench for the fiber optic cable at this location.

Segment 2 (ANF to Sylmar) Report for 11/5/98: 2A (Copco Rd to Templin Hwy): Revegetation crews began work on the Liebre, EMP 42.27, and Forest Inn, 47.35, testing cutout sites. Activities were limited to low fire risk tasks. 2C (Templin Hwy to Castaic): Work continued on the block wall on the ANF access road to Whitaker Reduction Station. Revegetation work occurred at several locations between EMP 53.00 and 55.46. Welding, x-ray, coating, trenchbreaker building, and shading took place between EMP 55.46 and 56.06. All construction activities on Segment 2C within Los Angeles County Fire Department jurisdiction (south of EMP 52.55) were halted for the day at 10:30 AM due to high fire risk. 2B (Templin Hwy to Sylmar): F/O cable was spliced at the intersection of Live Oak Rd and The Old Road. The directional drill at Feedmill Rd continued to be set up. A F/O splice box was installed at Lyons check valve north of Calgrove. At the Calgrove bore, both risers were welded, x-rayed, and coated. Also F/O conduit was pulled through the bore. Just north of Sierra Hwy, 400 feet of roadway asphalt was ground in preparation for trenching. 100 feet of asphalt was applied to the backfilled trench north of Sunshine Cyn at EMP 73.19. 300 feet of slurry was poured in the trench under the I-5 overpass at Sunshine Cyn. F/O conduit was laid at that location, also. Pipe was lowered-in and welded south of Sunshine Cyn.

Segment 3A, 3B, & 3C (Sylmar to Los Angeles): 3A: Backfill neared completion of their work at Verdugo Wash today. Electrical work was done at the block valve just north of the Burbank Animal Shelter. Fiber optic cable was pulled between the Metrolink Station and north of the animal shelter. Ditch work continued at Hollywood Way Bridge tie in. NIGHT work is scheduled for Hollywood Way Bridge hanging. 3B: Crews continued to hang razor wire on Downey Bridge. 3C: A six inch directional bore was drilled under Roxford Street for the fiber optic conduit today. Ditch, welding, x-ray, and coating crews worked on the mainline near Sepulveda Boulevard. The remaining walls for the California Aqueduct were poured today. Backfill crews continued closing up the ditch south of Grapevine Creek.

Segment 4B (El Segundo to Lynwood) 11-5-98: Backfilling activities continued at 108 Th St. ESEMP 1.16, Santa Ana ESEMP 0.04, and at Alameda and Fernwood AV ESEMP 0.02. Construction activities continued at El Segundo station ESEMP 11.00. This crew was working on getting the soil out of the casing. The crew at Lynwood station EMP 0.0. was installing pips today.

Segment 5 (Lynwood to Long Beach) Report for 5 November 1998: With the pilot hole being successfully drilled yesterday, reaming began today. The bore crew at Wilmington encountered suspicious soils that are being set aside awaiting test results. Two bores are in progress.

Inspection Summary

In the past 24 hours, 4 inspection event(s) were reported for 3 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page